OpenStudy (s):

lim x->0 1/5(sin(5x)/ sin(x) )^(1/2) ?

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{1}{5}\sqrt{\frac{\sin 5x}{\sin x}}~~?\]